15 liter of solution contain 40% of sulphuric acid it mixed with 25 liter of solution containing 60% of acid Find the percentage of water in that solution
Found 2 solutions by Alan3354, JoelSchwartz:
Answer by Alan3354(69443)

15 liter of solution contain 40% of sulphuric acid it mixed with 25 liter of solution containing 60% of acid Find the percentage of water in tat solution
-----------------------
15*40 + 25*60 = 40*C (C = %age)

Answer by JoelSchwartz(130)

.6*15+25*.4
9+10=19 liters of water that is in 40 liters of solution
x=rate of water in solution
40x=19
x=19/40
x=.475
40*.475=19
The percentage is forty seven point five percent.